OpenStudy (anonymous):

How would you verify [tan^2(x)-1]/[sec^2(x)] = [tan(x)-cot(x)]/[tan(x)+cot(x)] ? Any help is appreciated, thank you!

OpenStudy (anonymous):

well start with the right side first we got tan=1/cot

OpenStudy (anonymous):

so (tan-cot)/(tan+cot) =(tan-1/tan)/(tan+1/tan) =((tan^2-1)/tan)/((tan^2+1)/tan) and then....im sure u know what to do :)

OpenStudy (anonymous):

Thank you so much! I really appreciate it. :)

OpenStudy (anonymous):

ur very welcome